A student heats up a liquid from an unknown initial temperature (T1) to 54 °C with the thermal energy of 6 575 cal. 

If the mass of the liquid is 4 kg and its specific heat capacity is 100 J/ kg- °C, (1 cal = 4.184 J; 1000 g = 1 kg)


1
Expert's answer
2021-09-06T02:33:48-0400

change in temperature = (100 - 25) = 75.0°C.

change in thermal energy = mass × specific heat capacity × change in temperature.

= 0.200 × 4,180 × 75.0.

= 62,700 J (62.7kJ)